Make a sketch if you need to and solve this problem using similar triangles.
A telephone pole has a shadow of 16 ft. when a person 5 ft. tall has a shadow of 6 ft. How tall is the pole?
Write answer as a mixed number.

Answer by Alan3354(69443)

A telephone pole has a shadow of 16 ft. when a person 5 ft. tall has a shadow of 6 ft. How tall is the pole?
-----------------
16/h = 6/5
6h = 80
h = 40/3 feet